6. A tank is in the form of a right circular cylinder with hemispherical ends. The overall length of the tank is

4 meters and the diameter of the hemisphere is 1 meter. If a pump discharges a fluid whose density is

1.2 kg m /Liter in this tank at a rate of225 liters per minute determine a.) Weight of liquid inside the tank if

it is half full. b) total time to fill the tank assuming it is initially empty.


1
Expert's answer
2021-09-09T00:33:20-0400

Let "d=" the diameter of the hemisphere, "l="the overall length of the tank

Find the volume of the hemisphere


"V_1=\dfrac{1}{2}\big(\dfrac{4}{3}\pi(\dfrac{d}{2})^3\big)=\dfrac{\pi d^3}{12}"

Find the volume of the right circular cylinder


"V_2=\pi\big(\dfrac{d}{2}\big)^2(l-2(\dfrac{d}{2}))=\dfrac{\pi d^2(l-d)}{8}"

Find the mass of liquid inside the tank if it is half full


"m=\rho(V_1+V_2)=\rho\dfrac{\pi d^2(2d+3l-3d)}{24}"

"=\rho\dfrac{\pi d^2(3l-d)}{24}"

"m=1.2\ \dfrac{kgm}{10^{-3}\ m^3}\cdot\dfrac{\pi (1\ m)^2(3(4\ m)-1\ m)}{24}"

"=550\pi\ kgm\approx1728\ kg"

Find the mass of liquid inside the tank if it is half full


"W=mg"

"W=550\pi\ kgm(9.81\ m/s^2)\approx16950\ N"


"V=2(\dfrac{1}{2})\big(\dfrac{4}{3}\pi(\dfrac{d}{2})^3\big)=\dfrac{\pi d^3}{6}"

b) Find the volume of the tank


"V=2V_1+V_2"

"=2(\dfrac{\pi d^3}{12})+\dfrac{\pi d^2(l-d)}{8}"

"=\dfrac{\pi d^2(4d+3l-3d)}{24}"

"=\dfrac{\pi d^2(3l+d)}{24}"

"V=\dfrac{\pi (1\ m)^2(3(4\ m)+1\ m)}{24}=\dfrac{13\pi}{24}\ m^3"

Find the total time to fill the tank assuming it is initially empty


"t=\dfrac{\dfrac{13\pi}{24}\ m^3}{0.225 \ m^3/min}\approx7.6\ min"